# Unveiling the Nature and Fate of the Almost-Dark Cloud AGC 226178 through Mapping

### v1

@@ Abstract
Aims. We aim to investigate the nature and evolutionary history of the ”almost dark” cloud AGC 226178, projected within the Virgo cluster, which exhibits an -to-stellar mass ratio of approximately 1000.
@@
Results. FAST data reveal a short, low-velocity tail extending towards the dwarf galaxy VCC 2034, previously proposed as a possible origin for AGC 226178. However, VCC 2034 itself exhibits a line-of-sight asymmetric feature and a cometary morphology that indicates a stripping event unrelated to AGC 226178. VLA observations reveal a velocity gradient across AGC 226178, along with a clumpy internal structure. The velocity dispersion within the cloud exceeds the thermal linewidth, indicating the presence of turbulence and/or unresolved random motions. The whole cloud cannot be gravitationally bound by the observed atomic gas alone. The resolved clumps follow the standard mass–star formation rate relation and a tight mass–size relation, with those associated with star formation reaching surface densities above the theoretical threshold for self-shielding.
@@
Over the years, several objects have been detected which have been proposed as optically ‘dark galaxy’ candidates. With recent blind surveys, such as the Arecibo Legacy Fast ALFA blind survey (ALFALFA, Giovanelli et al. 2005), the FAST all sky HI survey (FASHI, Zhang et al. 2024), the Widefield ASKAP L-band Legacy All-sky Blind surveY(WALLABY, Koribalski et al. 2020), gas-rich objects with extremely faint or even no optical counterpart have been revealed. According to cosmological simulations in the prevalent paradigm, the expected number of dark matter subhalos around (say) a Milky Way size host galaxy is more than an order of magnitude larger than the detected satellite galaxies, which is known as the ‘missing satellite’ problem (Klypin et al. 1999). Either the standard CDM model is fundamentally flawed, or more low-mass halos would remain starless or extremely inefficient in star formation and thus still awaiting detection. In addition, different dark matter models predict different subhalo number distribution in the low-mass end (e.g. warm dark matter models; Bode et al. 2001). Furthermore, small dark halos with inefficient or no star formation and thus weak baryonic feedback effect are expected to retain the original dark matter halo structure. Therefore, identifying (almost) optically dark dwarf galaxies from observations may lead to crucial tests for both dark matter and galaxy formation models.
@@
Dark matter halos at the lower mass end are more prone to baryon depletion, which means they could be hard to observe. Below a characteristic halo mass, galaxy star formation may be completely suppressed due to severe baryon depletion and inefficient gas cooling (Hoeft et al. 2006) which is partly attributed to photoheating by the cosmological UV background. Hoeft & Gottlöber (2010) find the characteristic halo mass to be , which is robust against the assumed UV background flux density and resolution effects in their simulations. In addition, stellar feedback could also remove the gas component from galaxies with shallow potential wells. While it is generally challenging to detect stellar light from almost dark galaxies, it is relatively easier to detect them through neutral gas emission (e.g. Davies et al. 2006).
@@
Several candidates for (almost) dark galaxies have been reported in the past few decades, e.g., VIRGOHI21 (Davies et al. 2004), Nube (Montes et al. 2024), FAST J1039+4328 (Xu et al. 2023), and H1032-2819 (O’Beirne et al. 2024). To mention a few of the candidate dark galaxies, FAST J1039+4328 (Xu et al. 2023) shows a double-horned velocity profile, which is in line with a disc in regular rotation. Janowiecki et al. (2015) reported high-resolution observations of the isolated (no known neighbouring galaxies closer than ) system HI1232+20 (which consists of three almost dark clouds AGC 229383, AGC 229384, and AGC 229385). The main part of the system has a faint optical counterpart with a band peak surface brightness of . The two fainter gas clouds have no optical counterpart down to 27.9 and 27.8 mag in the band (Janowiecki et al. 2015). The nature of the system is still controversial, it lacks tidal features and a clear connection to other galaxies in the local environment.
@@
VIRGOHI21 was originally postulated to be a true dark galaxy (Minchin et al. 2005, 2007), however, subsequent observations (Haynes et al. 2007) revealed its elongated nature and faint gas connection with a larger galaxy. Simulations (Duc & Bournaud 2008) have also shown that this source is likely a tidal remnant. The dark cloud AGC 229101 has no optical counterpart in the SDSS imaging survey, but deeper observations (Leisman et al. 2021) revealed a faint stellar component. Higher resolution mapping indicates that it may be explained by a tidal encounter with neighbouring objects and a merger of two dark clouds.
@@
The origin of most of the observed dark galaxy candidates is still shrouded in mystery. Although highly elongated structures are better explained as remnants of tidal interactions (e.g., Duc & Bournaud 2008; Zhu et al. 2021; Yu et al. 2024) or ram pressure stripping (e.g., Serra et al. 2024), the origin of apparently roundish and isolated clouds is challenging our understanding: they genuinely could be dark galaxies. Beyond their potential importance in understanding the nature of dark matter, (almost) dark galaxy candidates are also valuable for studying the star formation process in an extremely -rich environment.
@@
With deep optical observations from the Next Generation Virgo cluster Survey (NGVS, Ferrarese et al. 2012), Junais et al. (2021) estimated the stellar mass of AGC 226178 as . They brought up the idea that AGC 226178 was resulted from ram pressure acting on the nearby Ultra-diffuse galaxy (UDG) NGVS 3543. This scenario might explain the faint bridge/clump of UV emission projected between the two objects (and also to the south of AGC 226178). Nevertheless, AGC 226178 and NGVS 3543 were later resolved into individual stars in Hubble Space Telescope (HST) imaging presented in Jones et al. (2022a). These authors found that NGVS 3543 is likely a foreground dwarf galaxy at a distance of 10 Mpc. Moreover, the metallicity expected for the mass of NGVS 3543 is more than an order of magnitude smaller than the measured nebular gas metallicity of AGC 226178. A re-analysis of the ALFALFA data by Jones et al. (2022a) revealed an extremely faint, bridge-like structure between AGC 226178 and the dwarf irregular galaxy VCC 2034. Considering the cluster environment in which both objects reside, they concluded that AGC 226178 likely originated from VCC 2034 through a ram pressure stripping event.
The origin of AGC 226178 remains inconclusive. Deeper and higher-resolution mapping of the gas is necessary to help resolve the ambiguity. AGC 226178 was recently observed by the Atomic gas in the Virgo Interacting Dwarf galaxies (AVID, Zhang et al. in prep) project, which obtained both multi-configuration VLA observations and deep FAST mapping of the emission line for a sample of dwarf galaxies involved in tidal interactions or dwarf-dwarf galaxy mergers. AGC 226178 was not among the core sample of AVID, but it falls into one of the AVID fields. In this paper, we perform a joint analysis of the AVID observations and ancillary multi-frequency data of AGC 226178, in order to shed further light on the origin and evolution of this intriguing, almost dark object.
@@ 2.1 VLA data
In this work, we utilize 21-cm emission line observations of AGC 226178 obtained with the VLA. The data consist of B, C and D multi-configuration observations from the AVID project (PID: 21A-231; PI: Hong-Xin Zhang), and the archival D-configuration data first presented by Cannon et al. (2015).
@@
The VLA data calibration and imaging were performed with the Common Astronomy Software Applications CASA
(McMullin et al. 2007) version 6.5.4-9. Particularly, the flux and phase calibration was performed primarily through the VLA Calibration Pipeline (adapted for emission line), with additional extensive manual flagging of data affected by radio frequency interference (RFI) by using the PLOTMS
and FLAGDATA
tasks in CASA. Hanning smoothing was applied to the data to yield an effective spectral resolution of . We performed continuum subtraction using the task UVCONTSUB
. Each execution block was calibrated and continuum-subtracted separately.
@@
All of the above calibrated and continuum-subtracted multi-configuration datasets were combined and imaged with tclean
. Two weighting methods were used for imaging. One method is called natural weighting and the other is Briggs’ weighting with a robust parameter set at =0.5 (hereafter robust-weighted). The natural weighting produces an image with the lowest noise, but at the expense of somewhat lower angular (and thus spatial) resolution, whereas the robust weighting attempts to balance the sensitivity and resolution. In addition to the above standard weighting schemes, we also applied a taper of
@@
To produce the imaging cubes, we run the tclean
task for each weighting scheme, applying a circular clean box centered on the source with a radius of . The deconvolution was carried out down to a - noise level.
The resulting beam size is for the robust-weighted image cube, for the natural-weighted image cube and for the -tapered natural weighted cube. The resulting rms noise is , , and for the robust-weighted, natural-weighted, and -tapered image cubes, respectively. This corresponds to a minimum detectable column density of , , and (3 over ), respectively.
The final source masks used for generating moment maps are also generated by SoFiA 2
, with the S+C spatial smoothing kernels set to 0 and 1.5 times the synthesised beam size, the spectral smoothing kernels to 0
and 3 channels, and a detection threshold of
3- over each smoothing scale. We limit the first moment (intensity-weighted mean velocity field) and second moment (velocity dispersion map) maps to only include the pixels where the signal-to-noise ratio is larger than 3.
@@ 2.2 FAST data
We obtained deep observations of AGC 226178 with FAST as part of the AVID project (PID: PT2022-0005; PI: Hong-Xin Zhang). FAST is the largest single-dish radio telescope in the world, with a 500 m diameter aperture (300 m illuminated aperture) (Nan et al. 2011; Jiang et al. 2019; Qian et al. 2020). Its L-band 19-beam receivers have a frequency range of 1.05-1.45 GHz in four polarisation modes and the half-power beam width (HPBW) of each beam is 2.90′ at 1.4 GHz (3.24′ after gridding in imaging process). We used the Multibeam on-the-fly (OTF) mode to uniformly map a 30′ 30′ sky area centred on VCC2037. AGC 226178 is about 12′ to the north-east of the field centre. Every observing block consists of a pair of horizontal and vertical scans, with a 23.4∘/53.4∘ (horizontal/vertical) rotation of the receiver, a 21.66′ scan separation, and 15 scan speed. We followed the observing strategy described in Wang et al. (2023) and Yang et al. (2025). The Spec(W+N) backend is used to record the data with a sampling time of 0.5 s, covering 500 MHz bandwidth with 65536 channels. The frequency and velocity resolution are 7.63 kHz and 1.61 km s-1 at 1.42 GHz. The flux calibration is accompanied by injecting a 10 K noise diode for 1s every 60s on target in order to calibrate the antenna temperature.
The observations were carried out on August 13, 21, 23, and September 4 of 2022. The accumulated on-source time per sky position is about 4.87 minutes. The raw data were reduced mainly using the FAST pipeline222https://hifast.readthedocs.io/en/v1.3/ developed by Jing et al. (2024). The pipeline follows the general reduction procedure for single-dish data. With FAST, we first choose the FFT-filter approach to remove standing waves (more details can be found in Xu et al. 2025). Then we use the ‘MedMed’ algorithm with the ‘poly-asym2’ method to perform the post-processing for baseline subtraction. After performing RFI flagging, frame conversion, and velocity correction, the final calibrated spectra are gridded into a data cube with a pixel size of 1′, The data cube is further Hanning smoothed to a velocity resolution of 6.4 km s-1. The final FAST data cube has an rms noise level of 0.78 prior to Hanning smoothing, which corresponds to a minimum detectable column density of (3 over 20 ).
As with the VLA data, we use the S+C algorithm in SoFiA 2
to search for genuine emission signal in the FAST image cube. Particularly, we choose spatial smoothing kernels of 0, 1 and 2 times the beam size, spectral smoothing kernels of 0, 5 and 12 , and a detection threshold of for each smoothing scale. We search for emission over a Heliocentric velocity range of .
@@ 3.1 Large-scale distribution
Fig.1 shows the FAST column density contours of the AGC 226178 field, overlaid on the DECaLS RGB image (Dey et al. 2019). The 3 sensitivity for a 20 line width of our FAST map is , which is approximately one-sixth of the detection limit of ALFALFA observations (, Jones et al. 2022a). Such a substantial improvement in sensitivity is expected to significantly enhance the ability to detect faint features on large scales. Complementarily, Fig. 2 displays the natural-weighted moments 0 contours from VLA observations, overlaid on the NGVS band image, providing the highest spatial resolution available for this field.
@@
The dwarf galaxies VCC 2037 and VCC 2034 can not be spatially distinguished in the FAST integrated column density map, due to their small angular separation. But they are well-separated in velocity, with a difference of in systemic velocity. AGC 226178 appears more extended towards the direction of VCC 2034 than in the opposite direction. As indicated in the position-velocity (PV) diagrams along the AGC 226178-VCC 2034 direction (dashed line in the left panel) shown in the right panel of Fig.1, AGC 226178 exhibits a velocity gradient along the direction, particularly with a faint, extended component towards the lower-velocity end. The overall morphology of AGC 226178 from FAST shows a compact main body and an extended, kinematically distinct tail component with central radial velocity of 1550 km s-1, consistent with a head-tail structure expected from ongoing stripping. The two components in the moment-0 map are integrated over the velocity ranges 1535–1565 for the tail, and 1565–1625 for the main body (see middle panel of Fig.1). The angular separation between the flux centres of main body and tail is measured to be 2.24 arcminutes. Assuming the Virgo cluster distance, this corresponds to a projected physical separation of 10.7 kpc.
@@
In a recent study, Dey et al. (2025) discovered another Blue Cloud (BC12) near AGC 226178 (referred to as BC3 in their paper), at a projected distance of only 9 kpc. The location of BC12 is marked with a blue star in the middle panel of Fig. 1. Observations of the H emission line indicate that the velocity of BC12 is consistent with that of AGC 226178, within the limits of the spectral resolution and observational uncertainties. The two systems also exhibit comparable metallicities. Moreover, the stellar component of BC12 appears to lie along the extension of the elongated optical structure of BC3. Therefore, it is highly likely that there was a physical connection between these two objects. Nevertheless, the position of BC12 is clearly offset from the direction of the tail detected in our FAST observations. BC12 is not detected in our VLA images, we therefore place a 3 upper limit on the neutral content of BC12 of , by adopting a 15″-diameter circular aperture and a 10.2 km/s linewidth for the tapered natural-weighted image cube.
@@
We note that, Jones et al. (2022a) found a low signal-to-noise connection between AGC 226178 and VCC 2034, based on the ALFALFA observation. Our deeper FAST observation reveals a short extension towards VCC 2034, but disproves a direct connection between the two. Moreover, the right PV diagram in Fig.1 also reveals an asymmetric extension in the line-of-sight velocity distribution towards the lower velocity end of VCC 2034. In contrast to the lower-velocity end, the contours at the higher-velocity end of VCC 2034 appear relatively squeezed, which is an indication of ongoing stripping. The stripping direction in velocity space points away from, rather than toward, AGC 226178. Although VCC 2034 is not spatially resolved in the FAST observation, the VLA image reveals a cometary morphology, with the head oriented roughly towards the north, instead of along the direction to AGC 226178 (Figs. 2 and 3). Therefore, both the morphology and the radial velocity—along with the direction of its gradient—of the stripping features argue against the possibility that AGC 226178 originated from VCC 2034. The misalignment of the stripping tails in both systems with respect to the direction towards the Virgo cluster centre suggests that they possess significant tangential motion within the cluster, rather than undergoing purely radial infall.
@@
The angular separation between AGC 226178 and its low surface brightness neighbour NGVS 3543 is only 1 arcmin, hence they cannot be spatially distinguished in the FAST data. NGVS 3543 does not yet have an detection, but according to the model predictions by Junais et al. (2022), if AGC 226178 originated from a ram pressure stripping event of NGVS 3543, then NGVS 3543 should still possess gas with a mass of . Therefore, we attempted to search for such a faint source using our data. Junais et al. (2022) estimated the stellar mass of NGVS 3543 . According to the Baryonic Tully-Fisher relation (BTFR) from Papastergis et al. (2016), the possible maximum rotation velocity of NGVS 3543 is . Assuming the full width at 50% of the peak intensity (Ponomareva et al. 2021), we expect to detect an mass down to with a significance of . If NGVS 3543 contains gas as expected by Junais et al. (2022), then it would be near the detection limit of the FAST observation. According to the Systematically Measuring Ultra-diffuse Galaxies (SMUDGes) survey (Zaritsky et al. 2023), the approximate velocity of NGVS3543 is reported as , estimated using the Distance-by-Association method. This approach assigns a redshift to ultra-diffuse galaxies based on the median redshift of at least three bright galaxies located within a projected distance of 1.5 Mpc. However, such estimates are highly uncertain for nearby systems, particularly within the Virgo cluster. While we took this velocity into consideration, we also searched for emission over a broader velocity range.
@@
To probe any emission from NGVS 3543 based on the FAST data, we used an aperture of one average beam size in diameter centred on the optical centre of NGVS 3543 (red dash circle in the left panel of Fig.1) to extract the spectrum. The spectrum is shown as the red dashed curve in Fig.4. Due to the limited spatial resolution of the FAST data, the aperture also includes emission from the main body of AGC 226178. No significant signal is found near the expected velocity of NGVS 3543. Given the notoriously unreliable nature of projected positions near clusters—particularly in the unrelaxed Virgo cluster—the galaxy’s true velocity could plausibly fall anywhere near or below 3000 . For this reason, we extended our search to a broader velocity range (500–4000 ), but still detected no emission, unless NGVS 3543’s velocity were to coincide with that of AGC 226178 and create a blended emission feature.
@@
We applied the same aperture-based search to the natural-weighted VLA data cube. No significant detection was found within a beam-sized aperture centred on NGVS 3543. Therefore, we did not find evidence for emission associated with NGVS 3543. Based on the noise level of the data, we place a 3 upper limit of on its mass.
Galaxy ID | Flux | SNR | Distance | ||||

### v2

@@ Abstract
Aims. We investigated the nature and fate of the almost dark cloud AGC 226178, projected within the Virgo cluster, which exhibits an -to-stellar mass ratio of approximately 1000.
@@
Results. The FAST data reveal a short low-velocity tail extending towards the dwarf galaxy VCC 2034, which was previously proposed as a possible origin for AGC 226178. However, VCC 2034 itself exhibits a line-of-sight asymmetric feature and a cometary morphology that indicates a stripping event unrelated to AGC 226178. VLA observations reveal a velocity gradient across AGC 226178, along with a clumpy internal structure. The velocity dispersion within the cloud exceeds the thermal line width, indicating the presence of turbulence and/or unresolved random motions. The whole cloud cannot be gravitationally bound by the observed atomic gas alone. The resolved clumps follow the standard mass–star formation rate relation and a tight mass–size relation: those associated with star formation reach surface densities above the theoretical threshold for self-shielding.
@@
Over the years, several objects have been detected which have been proposed as optically dark galaxy candidates. With recent blind surveys, such as the Arecibo Legacy Fast ALFA blind survey (ALFALFA, Giovanelli et al. 2005), the FAST all sky HI survey (FASHI, Zhang et al. 2024), and the Widefield ASKAP L-band Legacy All-sky Blind surveY(WALLABY, Koribalski et al. 2020), gas-rich objects with extremely faint or even no optical counterpart have been revealed. According to cosmological simulations in the prevalent paradigm, the expected number of dark matter subhalos around a Milky Way-size host galaxy is more than an order of magnitude larger than the detected satellite galaxies, which is known as the ‘missing satellite’ problem (Klypin et al. 1999). Either the standard CDM model is fundamentally flawed, or more low-mass halos would remain starless or extremely inefficient in star formation and thus still awaiting detection. In addition, different dark matter models predict different subhalo number distribution in the low-mass end (e.g. warm dark matter models; Bode et al. 2001). Furthermore, small dark halos with inefficient or no star formation, and thus weak baryonic feedback effects, are expected to retain the original dark matter halo structure. Therefore, identifying (almost) optically dark dwarf galaxies from observations may lead to crucial tests for both dark matter and galaxy formation models.
@@
Dark matter halos at the lower-mass end are more prone to baryon depletion, which means they could be hard to observe. Below a characteristic halo mass, galaxy star formation may be completely suppressed due to severe baryon depletion and inefficient gas cooling (Hoeft et al. 2006), which is partly attributed to photoheating by the cosmological ultraviolet (UV) background. Hoeft & Gottlöber (2010) find the characteristic halo mass to be , which is robust against the assumed UV background flux density and resolution effects in their simulations. In addition, stellar feedback could also remove the gas component from galaxies with shallow potential wells. While it is generally challenging to detect stellar light from almost dark galaxies, it is easier to detect them through neutral gas emission (e.g. Davies et al. 2006).
@@
Several candidates for (almost) dark galaxies have been reported in the past few decades: VIRGOHI21 (Davies et al. 2004), Nube (Montes et al. 2024), FAST J1039+4328 (Xu et al. 2023), and H1032-2819 (O’Beirne et al. 2024). To mention a few of the candidate dark galaxies, FAST J1039+4328 (Xu et al. 2023) shows a double-horned velocity profile, which is in line with a disc in regular rotation. Janowiecki et al. (2015) reported high-resolution observations of the isolated (no known neighbouring galaxies closer than ) system HI1232+20 (which consists of three almost dark clouds: AGC 229383, AGC 229384, and AGC 229385). The main part of the system has a faint optical counterpart with a -band peak surface brightness of . The two fainter gas clouds have no optical counterpart down to 27.9 and 27.8 mag in the -band (Janowiecki et al. 2015). The nature of the system is still controversial; it lacks tidal features and a clear connection to other galaxies in the local environment.
@@
VIRGOHI21 was originally postulated to be a true dark galaxy (Minchin et al. 2005, 2007); however, subsequent observations (Haynes et al. 2007) revealed its elongated nature and faint gas connection with a larger galaxy. Simulations (Duc & Bournaud 2008) have also shown that this source is likely a tidal remnant. The dark cloud AGC 229101 has no optical counterpart in the SDSS imaging survey, but deeper observations (Leisman et al. 2021) revealed a faint stellar component. Higher-resolution mapping indicates that it may be explained by a tidal encounter with neighbouring objects and a merger of two dark clouds.
@@
The origin of most of the observed dark galaxy candidates is still shrouded in mystery. Although highly elongated structures are better explained as remnants of tidal interactions (e.g. Duc & Bournaud 2008; Zhu et al. 2021; Yu et al. 2024) or ram pressure stripping (e.g. Serra et al. 2024), the origin of apparently roundish and isolated clouds is challenging our understanding: they genuinely could be dark galaxies. Beyond their potential importance in understanding the nature of dark matter, (almost) dark galaxy candidates are also valuable for studying the star formation process in an extremely -rich environment.
@@
With deep optical observations from the Next Generation Virgo cluster Survey (NGVS, Ferrarese et al. 2012), Junais et al. (2021) estimated the stellar mass of AGC 226178 as . They brought up the idea that AGC 226178 resulted from ram pressure acting on the nearby ultra-diffuse galaxy (UDG) NGVS 3543. This scenario might explain the faint bridge or clump of UV emission projected between the two objects (and also to the south of AGC 226178). Nevertheless, AGC 226178 and NGVS 3543 were later resolved into individual stars in Hubble Space Telescope (HST) imaging presented in Jones et al. (2022a). These authors found that NGVS 3543 is likely a foreground dwarf galaxy at a distance of 10 Mpc. Moreover, the metallicity expected for the mass of NGVS 3543 is more than an order of magnitude lower than the measured nebular gas metallicity of AGC 226178. A re-analysis of the ALFALFA data by Jones et al. (2022a) revealed an extremely faint, bridge-like structure between AGC 226178 and the dwarf irregular galaxy VCC 2034. Considering the cluster environment in which the two objects reside, they concluded that AGC 226178 likely originated from VCC 2034 through a ram pressure stripping event.
The origin of AGC 226178 remains inconclusive. Deeper and higher-resolution mapping of the gas is necessary to help resolve the ambiguity. AGC 226178 was recently observed by the Atomic gas in the Virgo Interacting Dwarf galaxies (AVID, Zhang et al. in prep.) project, which obtained both multi-configuration VLA observations and deep FAST mapping of the emission line for a sample of dwarf galaxies involved in tidal interactions or dwarf-dwarf galaxy mergers. AGC 226178 was not among the core sample of AVID, but it falls into one of the AVID fields. For this paper we performed a joint analysis of the AVID observations and ancillary multi-frequency data of AGC 226178 in order to shed further light on the origin and evolution of this intriguing, almost dark object.
@@ 2.1 VLA data
In this work, we use 21-cm emission line observations of AGC 226178 obtained with the VLA. The data consist of B, C and D multi-configuration observations from the AVID project (PID: 21A-231; PI: Hong-Xin Zhang), and the archival D-configuration data first presented by Cannon et al. (2015).
@@
The VLA data calibration and imaging were performed with the Common Astronomy Software Applications (CASA) (McMullin et al. 2007) version 6.5.4-9. Particularly, the flux and phase calibration was performed primarily through the VLA Calibration Pipeline (adapted for emission line), with additional extensive manual flagging of data affected by radio frequency interference (RFI) by using the PLOTMS and FLAGDATA tasks in CASA. Hanning smoothing was applied to the data to yield an effective spectral resolution of . We performed continuum subtraction using the task UVCONTSUB. Each execution block was calibrated and continuum-subtracted separately.
@@
All of the above calibrated and continuum-subtracted multi-configuration datasets were combined and imaged with tclean. Two weighting methods were used for imaging. One method is natural weighting; the other is Briggs’ weighting with a robust parameter set at =0.5 (hereafter robust weighting). The natural weighting produces an image with the lowest noise, but at the expense of somewhat lower angular (and thus spatial) resolution, whereas the robust weighting attempts to balance the sensitivity and resolution. In addition to the above standard weighting schemes, we also applied a taper of
@@
To produce the imaging cubes, we run the tclean task for each weighting scheme, applying a circular clean box centred on the source with a radius of . The deconvolution was carried out down to a - noise level.
The resulting beam size is for the robust-weighted image cube, for the natural-weighted image cube and for the -tapered natural weighted cube. The resulting root mean square (RMS) noise is , , and for the robust-weighted, natural-weighted, and -tapered image cubes, respectively. This corresponds to a minimum detectable column density of , , and (3 over ), respectively.
The final source masks used for generating moment maps are also generated by SoFiA 2, with the S+C spatial smoothing kernels set to 0 and 1.5 times the synthesized beam size, the spectral smoothing kernels to 0 and 3 channels, and a detection threshold of 3- over each smoothing scale. We limited the first moment (intensity-weighted mean velocity field) and second moment (velocity dispersion map) maps to only include the pixels where the signal-to-noise ratio is larger than 3.
@@ 2.2 FAST data
We obtained deep observations of AGC 226178 with FAST as part of the AVID project (PID: PT2022-0005; PI: Hong-Xin Zhang). FAST is the largest single-dish radio telescope in the world, with a 500 m diameter aperture (300 m illuminated aperture) (Nan et al. 2011; Jiang et al. 2019; Qian et al. 2020). Its L-band 19-beam receivers have a frequency range of 1.05-1.45 GHz in four polarization modes and the half-power beam width (HPBW) of each beam is 2.90′ at 1.4 GHz (3.24′ after gridding in imaging process). We used the Multibeam on-the-fly (OTF) mode to uniformly map a 30′ 30′ sky area centred on VCC2037. AGC 226178 is about 12′ to the north-east of the field centre. Every observing block consists of a pair of horizontal and vertical scans, with a 23.4∘/53.4∘ (horizontal/vertical) rotation of the receiver, a 21.66′ scan separation, and 15 scan speed. We followed the observing strategy described in Wang et al. (2023) and Yang et al. (2025). The Spec(W+N) backend was used to record the data with a sampling time of 0.5 s, covering 500 MHz bandwidth with 65536 channels. The frequency and velocity resolution are 7.63 kHz and 1.61 km s-1 at 1.42 GHz. The flux calibration was accompanied by injecting a 10 K noise diode for 1s every 60s on target in order to calibrate the antenna temperature.
The observations were carried out on August 13, 21, 23, and September 4 of 2022. The accumulated on-source time per sky position is about 4.87 minutes. The raw data were reduced mainly using the FAST pipeline222https://hifast.readthedocs.io/en/v1.3/ developed by Jing et al. (2024). The pipeline follows the general reduction procedure for single-dish data. With FAST, we first chose the FFT-filter approach to remove standing waves (more details can be found in Xu et al. 2025). Then we used the MedMed algorithm with the poly-asym2 method to perform the post-processing for baseline subtraction. After performing RFI flagging, frame conversion, and velocity correction, the final calibrated spectra were gridded into a data cube with a pixel size of 1′. The data cube was further Hanning smoothed to a velocity resolution of 6.4 km s-1. The final FAST data cube has an RMS noise level of 0.78 prior to Hanning smoothing, which corresponds to a minimum detectable column density of (3 over 20 ).
As with the VLA data, we used the S+C algorithm in SoFiA 2 to search for genuine emission signal in the FAST image cube. Particularly, we chose spatial smoothing kernels of 0, 1 and 2 times the beam size, spectral smoothing kernels of 0, 5 and 12 , and a detection threshold of for each smoothing scale. We searched for emission over a Heliocentric velocity range of .
@@ 3.1 Large-scale distribution
Fig.1 shows the FAST column density contours of the AGC 226178 field, overlaid on the Dark Energy Camera Legacy Survey (DECaLS) false-colour image (Dey et al. 2019). The 3 sensitivity for a 20 line width of our FAST map is , which is approximately one-sixth of the detection limit of ALFALFA observations (, Jones et al. 2022a). Such a substantial improvement in sensitivity is expected to significantly enhance the ability to detect faint features on large scales. Complementarily, Fig. 2 displays the natural-weighted moments 0 contours from VLA observations, overlaid on the NGVS -band image, providing the highest spatial resolution available for this field.
@@
The dwarf galaxies VCC 2037 and VCC 2034 cannot be spatially distinguished in the FAST integrated column density map, due to their small angular separation. But they are well-separated in velocity, with a difference of in systemic velocity. AGC 226178 appears more extended towards the direction of VCC 2034 than in the opposite direction. As indicated in the position-velocity (PV) diagrams along the AGC 226178-VCC 2034 direction (dashed line in the left panel) shown in the right panel of Fig.1, AGC 226178 exhibits a velocity gradient along the direction, particularly with a faint, extended component towards the lower-velocity end.
The overall morphology of AGC 226178 from FAST shows a compact main body and an extended, kinematically distinct tail component with central radial velocity of 1550 km s-1, consistent with a head-tail structure expected from ongoing stripping. The two components in the moment-0 map are integrated over the velocity ranges 1535–1565 for the tail, and 1565–1625 for the main body (see middle panel of Fig.1). The angular separation between the flux centres of main body and tail is measured to be 2.24 arcminutes. Assuming the Virgo cluster distance, this corresponds to a projected physical separation of 10.7 kpc.
@@
In a recent study, Dey et al. (2025) discovered another Blue Cloud (BC12) near AGC 226178 (referred to as BC3 in their paper), at a projected distance of only 9 kpc. The location of BC12 is marked with a blue star in the middle panel of Fig. 1. Observations of the H emission line indicate that the velocity of BC12 is consistent with that of AGC 226178, within the limits of the spectral resolution and observational uncertainties. The two systems also exhibit comparable metallicities. Moreover, the stellar component of BC12 appears to lie along the extension of the elongated optical structure of BC3. Therefore, it is highly likely that there was a physical connection between these two objects. Nevertheless, the position of BC12 is clearly offset from the direction of the tail detected in our FAST observations.
BC12 is not detected in our VLA images, we therefore place a 3 upper limit on the neutral content of BC12 of , by adopting a 15″-diameter circular aperture and a 10.2 km/s linewidth for the tapered natural-weighted image cube.
@@
We note that Jones et al. (2022a) found a low signal-to-noise connection between AGC 226178 and VCC 2034, based on the ALFALFA observation. Our deeper FAST observation reveals a short extension towards VCC 2034, but disproves a direct connection between the two. Moreover, the right PV diagram in Fig.1 also reveals an asymmetric extension in the line-of-sight velocity distribution towards the lower-velocity end of VCC 2034. In contrast to the lower-velocity end, the contours at the higher-velocity end of VCC 2034 appear relatively squeezed, which is an indication of ongoing stripping. The stripping direction in velocity space points away from, rather than towards, AGC 226178. Although VCC 2034 is not spatially resolved in the FAST observation, the VLA image reveals a cometary morphology, with the head orientated roughly towards the north, instead of along the direction to AGC 226178 (Figs. 2 and 3). Therefore, both the morphology and the radial velocity—along with the direction of its gradient—of the stripping features argue against the possibility that AGC 226178 originated from VCC 2034. The misalignment of the stripping tails in both systems with respect to the direction towards the Virgo cluster centre suggests that they possess significant tangential motion within the cluster, rather than undergoing purely radial infall.
@@
The angular separation between AGC 226178 and its low surface brightness neighbour NGVS 3543 is only 1 arcmin; hence, they cannot be spatially distinguished in the FAST data. NGVS 3543 does not yet have an detection, but according to the model predictions by Junais et al. (2022), if AGC 226178 originated from a ram pressure stripping event of NGVS 3543, then NGVS 3543 should still possess gas with a mass of . Therefore, we attempted to search for such a faint source using our data. Junais et al. (2022) estimated the stellar mass of NGVS 3543 . According to the Baryonic Tully-Fisher relation (BTFR) from Papastergis et al. (2016), the possible maximum rotation velocity of NGVS 3543 is . Assuming the full width at 50% of the peak intensity (Ponomareva et al. 2021), we expect to detect an mass down to with a significance of . If NGVS 3543 contains gas as expected by Junais et al. (2022), then it would be near the detection limit of the FAST observation. According to the Systematically Measuring Ultra-diffuse Galaxies (SMUDGes) survey (Zaritsky et al. 2023), the approximate velocity of NGVS3543 is reported as , estimated using the Distance-by-Association method. This approach assigns a redshift to ultra-diffuse galaxies based on the median redshift of at least three bright galaxies located within a projected distance of 1.5 Mpc. However, such estimates are highly uncertain for nearby systems, particularly within the Virgo cluster. While we took this velocity into consideration, we also searched for emission over a broader velocity range.
@@
To probe any emission from NGVS 3543 based on the FAST data, we used an aperture of one average beam size in diameter centred on the optical centre of NGVS 3543 (red dash circle in the left panel of Fig.1) to extract the spectrum. The spectrum is shown as the red dashed curve in Fig.4. Due to the limited spatial resolution of the FAST data, the aperture also includes emission from the main body of AGC 226178.
@@
No significant signal is found near the expected velocity of NGVS 3543. Given the notoriously unreliable nature of projected positions near clusters—particularly in the unrelaxed Virgo cluster—the galaxy’s true velocity could plausibly fall anywhere near or below 3000 . For this reason, we extended our search to a broader velocity range (500–4000 ), but still detected no emission, unless NGVS 3543’s velocity were to coincide with that of AGC 226178 and created a blended emission feature. We applied the same aperture-based search to the natural-weighted VLA data cube. No significant detection was found within a beam-sized aperture centred on NGVS 3543. Therefore, we did not find evidence for emission associated with NGVS 3543. Based on the noise level of the data, we place a 3 upper limit of on its mass.
| Galaxy ID | Flux | S/N | Distance | ||||